OpenStudy (anonymous):

integral sqrt(x^2-1)dx

OpenStudy (lgbasallote):

are you familiar with the method of trigonometric substitution/

OpenStudy (anonymous):

yes. i can get it to cot^2csc

OpenStudy (anonymous):

-cot^2csc**

OpenStudy (lgbasallote):

\[\int -\cot^2 \theta \csc \theta d \theta\] ??

OpenStudy (anonymous):

exactly. but from there idk where to go

OpenStudy (lgbasallote):

you can try and change it into sines and cosines first

OpenStudy (anonymous):

\[\frac{ \cos^2 }{ \sin^3 }\]